3 years ago Frank was 5 and his mother was 5 times his age. How old will his mother be in 4 years' time? How old will Frank be in 4 years' time?
Answer by partha_ban(41)

3 years age, mother's age = 5 * 5 = 25 years
Current age:
Frank = 5 + 3 = 8 years
Mother = 25 + 3 = 28 years
After 4 years:
Frank = 8 + 4 = 12 years
Mother = 28 + 4 = 32 years